Two factory plants are making TV panels. Yesterday, Plant A produced
4000
panels. Two percent of the panels from Plant A and
5%
of the panels from Plant B were defective. How many panels did Plant B produce, if the overall percentage of defective panels from the two plants was
4%
?

Answer by advanced_Learner(501)

0.02x+0.05y=0.04(x+y)
0.02x+0.05y=0.04(4000+y)
0.02(4000)+0.05y=0.04(4000+y)
y=8000